OpenStudy (anonymous):

I have gotten myself all confused now... (sin x)/(sin x)^2 +1....I can cancel the sin x? brain is saying yes other part of brain is saying are you sure? other part of brain is saying you will regret it....not sure who to listen to

OpenStudy (mertsj):

If you cancel it, you are assuming that it is not 0. Do you really want to make that assumption?

OpenStudy (mertsj):

Also is the problem:\[\frac{\sin x}{\sin^2x+1} or \frac{\sin x}{\sin^2x}+1\]

OpenStudy (anonymous):

the one is in the denominator so example 1

OpenStudy (mertsj):

Then absolutely not. It cannot be cancelled. Only factors may be cancelled. NOT addends.

OpenStudy (mertsj):

So what exactly are you supposed to be doing with this problem?

OpenStudy (anonymous):

I wrote myself into a corner trying to answer a question from purplemouse way down the page....I think my attempt hit a deadend...sadly lol

OpenStudy (mertsj):

Good luck

OpenStudy (anonymous):

∫1/((sinx)2+1)∗(sinx)/1 seems like there should be a way out of here...

OpenStudy (mertsj):

Where is Turing or Satellite when you need them

OpenStudy (anonymous):

hmm that should be an sinx^2 in the denominator...didnt paste well

OpenStudy (anonymous):

\[\frac{\sin(x)}{\sin^2(x)}=\sin(x)\] for sure

OpenStudy (anonymous):

but not \[\frac{\sin(x)}{\sin^2(x)+1}\]

OpenStudy (anonymous):

see there I should listen to my brain...just not sure which part

OpenStudy (mertsj):

\[\int\limits_{?}^{?}\frac{sinx}{\sin^2x+1}dx\]

OpenStudy (mertsj):

Isn't that the problem?

OpenStudy (anonymous):

indeed

OpenStudy (anonymous):

hmm maybe \[\frac{\sin(x)}{\sin^2(x)+1}=\frac{\sin(x)}{2-\cos^(x)}\] then a u - sub

OpenStudy (anonymous):

\[\frac{\sin(x)}{\sin^2(x)+1}=\frac{\sin(x)}{2-\cos^2(x)}\]

OpenStudy (mertsj):

u=cosx and du=-sin(x)dx

OpenStudy (anonymous):

\[\int\limits_{?}^{?}1/(u^2 + 2) du\]